Most of a cellMost of a cell's molecules can be classified as be's molecules can be classified as belonging to one of four major classes: lipids, carbohydrates, nucleotides, and amino acids. Sort the following molecules into one of the four major classes.
Answer #1
Concepts and reason

A list of compounds is given which needs to be identified as lipid, carbohydrates, nucleotide, and amino acids. Consider the general structural properties of the compounds and then classify the given structures into them.

Fundamentals

• Carbohydrates are polyhydroxy aldehydes or ketones which are composed of carbon, hydrogen, and oxygen having general formula CnH2nOn{{\rm{C}}_n}{{\rm{H}}_{{\rm{2}}n}}{{\rm{O}}_n} .

• Lipids are the naturally occurring organic compounds found in the living system. Fats, oils, and steroids belongs to lipid family.

• Amino acids contain an amine group (NH2)\left( {{\rm{N}}{{\rm{H}}_{\rm{2}}}} \right) , carboxylic acid group (COOH)\left( {{\rm{COOH}}} \right) , and an alkyl side chain (R)\left( {\rm{R}} \right) . Amino acids as the name suggests is formed from the carboxylic acid containing an amine group.

• Nucleotides are fundamental unit of nucleic acids. It contains a nitrogenous base, phosphoric acid, and a sugar residue.

Consider the given structures of the compound and identify the structure of lipid as shown below:

ҫH-ОН
о
но

The above structure represents a steroid thus, it is a lipid.

Consider the given structures of the compound and identify polyhydroxy ketones or aldehydes that represents the structure of carbohydrates as shown below:

Carbohydrate
Н
Н—с—Он
|С-H 00,
Но—с—н
н—с—он
Н—с—он
CH-ОН

Observe the given structures of the compound and identify the structure that contains a nitrogenous base, pentose sugar, and a phosphate group. The combination of these structure in a single compound represents the structure of nucleotide as shown below:

Nucleotide
NH2
N
-0—СH,
Н
Н
Н
ОН
Н
ОН
Z-
т.

Consider the given structures of the compound and identify the structure of amino acid as shown below:

Amino Acid
Н
Н,С—с—соо
-Соо
NH3

Ans:

Hence, the following molecules can be classified as lipid:

Lipid
ҫH,ОН
о
но

Hence, the following molecules can be classified as carbohydrates:

Carbohydrate
Н
Н—с—Он
|С-H 00,
Но—с—н
н—с—он
Н—с—он
CH-ОН

Hence, the following molecule can be classified as nucleotide:

Nucleotide
NH2
N
-0—СH,
Н
Н
Н
ОН
Н
ОН
Z-
т.

Hence, the following molecule can be classified as amino acid:

Amino Acid
Н
Н,С—с—соо
-Соо
NH3